OpenStudy (lovelyanna):

Which relation is a function? A. {(1, 2); (1, 3); (1, 4); (1, 5)} B. {(1, 2); (2, 3); (3, 4); (4, 5)} C. {(1, 2); (2, 3); (3, 4); (2, 5)} D. {(1, 2); (3, 2); (3, 3); (4, 2)} B?

OpenStudy (lovelyanna):

@jigglypuff314

OpenStudy (lovelyanna):

Emily.. how much time do you have to stay on?

jigglypuff314 (jigglypuff314):

B is correct :) and maybe 20 minutes...

OpenStudy (lovelyanna):

Ok, then I'll ask quickly!

OpenStudy (lovelyanna):

What is the range of the relation? {(1, 2), (2, 4), (3, 2), (4, 4)} A. {2, 4} B. {1, 2, 3, 4, 6} C. {2, 4, 6} D. {1, 2, 3, 4} A

jigglypuff314 (jigglypuff314):

correct :)

OpenStudy (lovelyanna):

Ok, next :) For the function f(x) = 2 – 3x, find f(–4). A. –5 B. 14 C. 12 D. –10

jigglypuff314 (jigglypuff314):

f(-4) means plug in -4 for x into the function f(x) so 2 - 3(-4) =

OpenStudy (lovelyanna):

B

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (lovelyanna):

Write the equation of the function graphed. A. f(x) = |x| – 4 B. f(x) = |x| + 4 C. f(x) = |x – 4| D. f(x) = |x + 4| http://static.k12.com/calms_media/media/242500_243000/242869/1/776b28b2042095efe0c0bbedfe0778ce32d29a11/VHS_ALG_L3_S2_01_16_UTQ4.gif

OpenStudy (lovelyanna):

C?

jigglypuff314 (jigglypuff314):

f(x) = |x - h| + k when (h, k) is the vertex D

OpenStudy (lovelyanna):

Oh.. so the answer is D?

jigglypuff314 (jigglypuff314):

yup yup :)

OpenStudy (lovelyanna):

thanks! on to the next one :) Which is the graph of the function g(x) = |x + 3|?

jigglypuff314 (jigglypuff314):

left 3 so D

OpenStudy (lovelyanna):

umm.. A?

OpenStudy (lovelyanna):

oh.. oops

OpenStudy (lovelyanna):

wait.. which question are you talking about?

jigglypuff314 (jigglypuff314):

g(x) = |x + 3|

OpenStudy (lovelyanna):

@jigglypuff314 its D for the graph?

OpenStudy (lovelyanna):

oh

jigglypuff314 (jigglypuff314):

D for both?

OpenStudy (lovelyanna):

D for this one: Which is the graph of the function g(x) = |x + 3|?

jigglypuff314 (jigglypuff314):

yup

OpenStudy (lovelyanna):

Which equation represents a direct linear variation? A. y = x2 B. 1/x C. 2/5x D. y = x + 3 C?

jigglypuff314 (jigglypuff314):

y = x is linear D

OpenStudy (lovelyanna):

is y=x+2 linear?

jigglypuff314 (jigglypuff314):

yup :)

OpenStudy (lovelyanna):

then why is it wrong in: Which equation represents a direct linear variation? A. y = ½x B. y = x + 2 C. 1/x D. y = x^2 A is correct

OpenStudy (lovelyanna):

C. y=1/x*

jigglypuff314 (jigglypuff314):

... ... ...? idk >,< :,(

OpenStudy (lovelyanna):

its ok. :) Which is the direct linear variation equation for the relationship? y varies directly with x and y = 12 when x = 4. A. y = x + 8 B. y = 2x + 4 C. y = x – 8 D. y = 3x D?

jigglypuff314 (jigglypuff314):

yeah, I think so :)

OpenStudy (lovelyanna):

Emily.. you sure all the others are correct?

jigglypuff314 (jigglypuff314):

I'm human, so ... 89% sure XD

OpenStudy (lovelyanna):

Which relationship shows a quadratic variation?

jigglypuff314 (jigglypuff314):

uhhhhh... between B or C most likely B 1, 4, 9, 16 pattern of quadratic...

OpenStudy (lovelyanna):

ok, you sure? =)

jigglypuff314 (jigglypuff314):

C...

jigglypuff314 (jigglypuff314):

yeah C... 1, 4, 9, 16...

OpenStudy (lovelyanna):

Which is the quadratic variation equation for the relationship? y varies directly with x2 and y = 27 when x = 6. A. y = x^2 + 25 B. y = 12x C. y = 4x^2 D. y = 2x^2

OpenStudy (lovelyanna):

So C, or B?

jigglypuff314 (jigglypuff314):

C for the other one... idk this one...

OpenStudy (lovelyanna):

you don't know this one at all..?

jigglypuff314 (jigglypuff314):

maybe like y = kx^2 ? but me not good with "direct variation" problems... sorry :/

OpenStudy (lovelyanna):

then C?

jigglypuff314 (jigglypuff314):

idk probably?

OpenStudy (lovelyanna):

ok :) Which relationship shows an inverse variation?

jigglypuff314 (jigglypuff314):

as x increases, y decreases... B

OpenStudy (lovelyanna):

huh? I think they are all supposed to equal the same

jigglypuff314 (jigglypuff314):

inverse variation is when x increases y decreases...

OpenStudy (lovelyanna):

its the first one...

jigglypuff314 (jigglypuff314):

decreases evenly...?

OpenStudy (lovelyanna):

which one is inverse here?

jigglypuff314 (jigglypuff314):

C

OpenStudy (lovelyanna):

nope, its D

OpenStudy (lovelyanna):

says so here

jigglypuff314 (jigglypuff314):

meh >,< me not good at variations...

OpenStudy (lovelyanna):

ok, so that one is A

jigglypuff314 (jigglypuff314):

suppose so :P you probably shouldn't ask me these questions any more XD

OpenStudy (lovelyanna):

y varies inversely with x and y = 8 when x = 2. Which is the inverse variation equation for the relationship? A. y = x + 2 B. y = 2x C. 16/2 D. y = 16x

OpenStudy (lovelyanna):

C?

OpenStudy (lovelyanna):

oh.. C is actually 16/x

jigglypuff314 (jigglypuff314):

y = k/x yeah yeah C

OpenStudy (lovelyanna):

ok, last one! Write a coordinate rule for the translation of f(x) to g(x). A. (x, y) → (x + 5, y – 4) B. (x, y) → (x + 4, y – 5) C.(x, y) → (x – 4, y + 5) D. (x, y) → (x – 5, y + 4)

OpenStudy (lovelyanna):

C..?

jigglypuff314 (jigglypuff314):

left 5 = (x-5,...)

OpenStudy (lovelyanna):

oh, D?

jigglypuff314 (jigglypuff314):

I think so. :)

OpenStudy (lovelyanna):

ok! thanks :)

jigglypuff314 (jigglypuff314):

you're better then me at this stuff at this point XD

OpenStudy (lovelyanna):

Lol, no, not really :P

jigglypuff314 (jigglypuff314):

at this stuff yes :P

OpenStudy (lovelyanna):

I'll be going now, thanks for your help! :)

jigglypuff314 (jigglypuff314):

see ya later! :)